# Systematic search for cool hypervelocity stars in multi-epoch WISE data

### v1

@@ 3 Discussion
Table 1 shows the full list of high proper motion objects detected in our analysis, along with their Simbad associations. We successfully recovered the fastest known brown dwarf, WISEA J085510.74-071442.5(Luhman, 2014) (upper rightmost point in Figure 4 – note that CatWISE2020 contains three different objects associated to it, as seen in right panel of Figure 3), as well as the fastest known star, Barnard’s Star (lower rightmost point), but did not detect anything moving faster than that, despite the algorithm specifically being tailored for catching the largest proper motions.
In order to assess the sensitivity of the algorithm for detecting other known cool objects with high proper motions we cross-matched our detections with the list of known cool dwarfs from UltraCoolSheet (Best et al., 2024, 2021, 2024) database. For that, we used CatWISE2020 positions from the database, and proceeded in the same way as we employed before. As it is seen Figure 5, most of the objects moving faster than 0.6′′/year are successfully detected in our analysis. We examined the data for several missing ones and found that they are either too faint to be detectable in epochal unTimely catalogues (in contrast e.g. to CatWISE2020 where detections are done in all-time coadded images which are significantly deeper), or situated in complicated fields with too few isolated detections.
3.1 Candidates without Simbad associations
Of all moving objects that we detected in our analysis, 42 are not listed in Simbad (Wenger et al., 2000) database, and thus may potentially be previously unknown brown dwarfs. To better assess them, we performed their association with several large catalogues both in infrared and optical range, by propagating their positions to the catalogue epoch where known. Specifically, we cross-matched the candidates with CatWISE2020 (Marocco et al., 2021), identified proper unique associations in ambiguous cases by comparing the brightness and proper motion (see Section 2.2 and right panel of Figure 3 for the discussion) and used their identifiers as a human-readable designation for the candidates from now on. We also cross-matched them with 2MASS point source catalogue (Skrutskie et al., 2006), the catalogue of Pan-STARRS DR2 epochal (“warp”) measurements (Chambers et al., 2016), VISTA Hemisphere Survey (McMahon et al., 2013), UKIRT Hemisphere Survey (Dye et al., 2018) epochal detections, SkyMapper DR4 (Onken et al., 2024) epochal detections, Sloan Digital Sky Survey DR16 (Ahumada et al., 2020), DENIS (Denis, 2005), Dark Energy Survey DR2 (Abbott et al., 2021), and DELVE DR2 (Drlica-Wagner et al., 2022). In cases of ambiguous matches (especially for Dark Energy Survey which is based on a stack of multi-epoch images with no precise times specified) we inspected individual candidates and selected the ones with most compatible brightness values. For multi-epoch Pan-STARRS DR2, SkyMapper DR4, and DELVE DR2 data, with numerous separate detections of every object at different epochs, we assumed the median value for their magnitudes in every filter.
Despite detection of at least four of our candidates in Dark Energy Survey (see Table 3), none of them is in the list of ultracool dwarf candidates from DES 6-year data (dal Ponte et al., 2023), which is consistent with relatively large proper motions of our candidates. On the other hand, we also performed the search for these candidates in the UltraCoolSheet (Best et al., 2024, 2021, 2024) compilative database of cool dwarfs555UltraCoolSheet is available online at http://bit.ly/UltracoolSheet., and found two of them – CWISE J234627.27+665732.6 (#11) and CWISE J035532.23+474358.1 (#33) – there, originally published by Tinney et al. (2018) and (Smith et al., 2014), correspondingly. One more – CWISE J124909.08+362116.0 (#4) – was recently published by Burgasser et al. (2024) as an L subdwarf at the star/brown dwarf mass limit. Seven more candidates are also listed in Brooks et al. (2024). It supports the possibility of the other ones also being suitable candidates for brown dwarfs.
The summary of the data we assembled on the candidates is given in Table 2, and the full photometry is given in Table 3. Using these measurements, we constructed two-color diagrams666For WISE colors here and below we use CatWISE2020 values for both W1 and W2 bands, instead of deriving these quantities from unTimely catalogue. It is motivated both by complexity of combining measurements from different unTimely epochs with varying quality and degree of crowding, as well as by supposedly better accuracy of CatWISE values derived by a combined fit to all available images at once (Marocco et al., 2021). The only exception for that is CWISE J105716.29-675728.4 (#36) – see below. that are shown in Figure 6, and overplotted the data for known cool dwarfs from Kirkpatrick et al. (2021) and UltraCoolSheet (Best et al., 2024) along with the template cool dwarfs colors from dal Ponte et al. (2023).
Right panel of Figure 6 shows that the spectral type for cool dwarfs is changing quite smoothly in the (K-W2) vs (J-W2) plane. We may exploit it to make a simple estimation of the spectral types for the candidates, by taking a median type value in their vicinity there. For the candidates lacking J and K magnitudes but having the color (W1-W2)¿0.5 we may use just it to also estimate the spectral type, albeit less reliably due to larger scatter. For (W1-W2)¡0.5 the dependence of the type on the color is too weak to estimate it. Table 2 shows these tentative types estimated from either colors. In both cases, the spread of catalogue values in the candidate vicinity defines the range of possible candidate types. For the candidates having multi-color data in optical and near-infrared bands (see Table 3 and Table 4) we also used the VO SED Analyzer (VOSA777VO SED Analyzer is available at http://svo2.cab.inta-csic.es/svo/theory/vosa/) (Bayo et al., 2008) to estimate the spectral type by template fitting, and got the results generally consistent with the ones we acquired from infrared colors.
In a similar manner, we may use the (W1-W2) vs absolute W2 magnitude plane (left panel of Figure 7) to estimate absolute magnitudes in W2 band for the candidates. To do so, for every candidate we selected all UltraCoolSheet objects falling within the 1 range of its (W1-W2) with known parallaxes, and computed the median and median absolute deviation of their absolute magnitudes, to be used as an estimates for the candidate absolute magnitude and its uncertainty, correspondingly. Then, using these values, we estimated the distances to the candidates that are shown in right panel of Figure 7, and listed in Table 2. The projected velocities corresponding to observed proper motions and estimate distances, are shown in Figure 8.
3.2 Individual objects
CWISE J105716.29-675728.4 (#36)
Association with this CatWISE2020 object is unreliable as it has inconsistent direction and amplitude of proper motion despite having consistent position and W2 magnitude. This is probably caused by a crowded field having several stationary objects around that confuse crowdsource pipeline. Therefore, we cannot fully believe in the magnitudes reported for this object in the catalogue, and thus computed them from unTimely epochal catalogues888For crowded fields like the one around CWISE J105716.29-675728.4, unTimely epochal magnitudes of individual objects often display prominent correlation of the magnitude with fracflux parameter describing the fraction of the flux inside PSF footpring coming from the object itself. To account for it, we performed a linear fit of this dependency, and extrapolated it to fracflux=1, i.e. fully uncrowded case, using the extrapolated value and its propagated error as magnitude estimation. As this procedure is not sufficiently stable, we did not perform similar fit for all other objects where CatWISE2020 associations are more reliable, using catalogue magnitudes instead. that give (W1-W2)=-0.080.06. Together with optical and near-infrared colors, it suggests that the object is an early K-type star, with VOSA best fit corresponding to K2 type. Therefore we excluded this object from the spectral type analysis and distance estimation.
CWISE J140808.52-491759.4 (#39)
has sufficiently large (W1-W2) to be classified as T5, but its near-infrared (J and K) magnitudes place it among M9-L0 types in right panel of Figure 6. However, there is a known spectroscopically classified T4.5 dwarf CWISE J020743.00+000056.1 (Geballe et al., 2002) close to its position, so we may assume that the classification based on (W1-W2) is correct for it.
CWISE J124909.08+362116.0 (#4)
has (W1-W2)=0.33 and so its type cannot be reliably estimated from that color alone. VOSA template fitting of its multi-wavelength data gives best-fit L1 spectral type which is consistent with that color, and with the location of the object in Figure 6 and Figure 7. It is also consistent with sdL1 type derived from the spectrum by Burgasser et al. (2024).
CWISE J234627.27+665732.6 (#9)
also has small (W1-W2)=0.2, but it lacks enough multi-wavelength measurements to achieve reasonable template fit with VOSA. Available data suggest late M or early L type.
CWISE J011128.00-103535.7 (#15)
has (W1-W2)=0.17, VOSA template fit gives L0 spectral type for it.
CWISE J204928.79-443146.9 (#14)
has T8-Y0 spectral type estimated from (W1-W2) color that is consistent with T7.3 type previously published by Tinney et al. (2018).
CWISE J035532.23+474358.1 (#40)
has T5-T7 spectral type estimated from (W1-W2) colors that is consistent with T6 type previously published by Smith et al. (2014).
J230252.77-131458.0 (#18)
is published in Brooks et al. (2024) with L1 spectral type at 14271 pc distance, but with very large of the fit. Our estimation is T4–T5 based on (W1-W2) color, and T4.5 based on VOSA fit, and the distance is 317 pc. Moreover, our direct estimation of proper motion is nearly twice as large as CatWISE2020 value employed by Brooks et al. (2024). Therefore, we tend to believe in our estimations, and consider the object reliable T-dwarf candidate.
J104222.34-834921.2 (#22)
is published in Brooks et al. (2024) with M8 spectral type, consistent with our estimation of M7–M9.
J110048.87-554006.8 (#24)
is published in Brooks et al. (2024) with M9 spectral type, consistent with our estimation of M6–L1.
J060302.31+321532.1 (#27)
is published in Brooks et al. (2024) with M8 spectral type, consistent with our estimation of M7–M9.
J194428.05+053344.8 (#31)
is published in Brooks et al. (2024) with M4 spectral type. Our estimation is M9–T3 based on (K-W2)-(J-W2) colors, and L6 based on VOSA fit.
J220134.02-041609.5 (#32)
is published in Brooks et al. (2024) with L1 spectral type, but with very large of the fit. Our estimation is T4–T6 based on (K-W2)-(J-W2) colors, and T6.5 based on VOSA fit. Thus, we again consider our estimation correct, and consider the object a reliable T dwarf candidate.
J093016.98+604631.2 (#41)
is published in Brooks et al. (2024) with T6 spectral type, consistent with our estimation of T5–T7. We consider it a reliable T dwarf candidate.
Reliable brown dwarf candidates.
Based on the positions in Figure 6 and Figure 7 and estimated spectral types we may conclude that at least the following 15 objects, in addition to previously published ones mentioned above, are also reliable late-type brown dwarf candidates that should be targets for more detailed investigation: CWISE J053046.20+440849.2 (#1), CWISE J062316.19+071505.6 (#3), CWISE J105018.24-683054.9 (#6), CWISE J214706.01-540407.6 (#7), CWISE J011816.64-723917.9 (#8), CWISE J140048.92+253759.8 (#10), CWISE J075744.48-300504.3 (#16), CWISE J045456.23-742522.6 (#21), CWISE J024641.43-634329.2 (#23), CWISE J095855.00+540305.4 (#25), CWISE J092710.37-474155.5 (#30), CWISE J102332.99+221527.6 (#33), CWISE J141403.72-171531.1 (#34), CWISE J184526.47+754829.7 (#35), CWISE J140808.52-491759.4 (#39), All of them have estimated distances closer than 50 pc and types later than T0.
| Unique ID | RA | Dec | Epoch | pmRA | pmDec | W2 | pm | pm Gaia | Simbad |
| MJD | arcsec/year | arcsec/year | mag | arcsec/year | arcsec/year | ||||
| 2688p045w2o0025372e001 | 269.4495 | 4.7242 | 55456 | -0.62 0.05 | 10.32 0.02 | 3.83 0.00 | 10.34 | 10.39 | Barnard’s star |
| 0788m455w2o0028019e000 | 77.9449 | -45.0345 | 55249 | 6.51 0.06 | -5.71 0.02 | 4.55 0.00 | 8.66 | – | HD 33793 |
| 1342m076w2o0020745e001 | 133.7953 | -7.2451 | 55321 | -8.10 0.05 | 0.66 0.02 | 13.92 0.04 | 8.13 | – | WISEA J085510.74-071442.5 |
| 1781p378w2o0011596e000 | 178.2595 | 37.7020 | 55344 | 4.00 0.01 | -5.88 0.03 | 4.08 0.00 | 7.11 | – | HD 103095 |
| 3469m364w2o0023991e000 | 346.4911 | -35.8491 | 55340 | 6.71 0.05 | 1.28 0.02 | 2.76 0.00 | 6.83 | – | HD 217987 |
| 0018m379w2o0026788e000 | 1.3724 | -37.3641 | 55352 | 5.62 0.04 | -2.33 0.01 | 3.96 0.00 | 6.09 | – | HD 225213 |
| 1225p090w2o0009243e000 | 122.9930 | 8.7585 | 55303 | 1.06 0.01 | -5.10 0.01 | 7.23 0.00 | 5.21 | 5.21 | Ross 619 |
| 0440p166w2o0017388e000 | 43.2638 | 16.8706 | 55228 | 3.42 0.02 | -3.79 0.01 | 7.16 0.00 | 5.10 | 5.12 | Teegarden’s Star b |
| 3174p393w2o0004832e000 | 316.7405 | 38.7589 | 55347 | 4.02 0.03 | 3.13 0.03 | 1.84 0.00 | 5.09 | – | * 61 Cyg A |
| 3174p393w2o0004657e000 | 316.7456 | 38.7514 | 55347 | 4.03 0.04 | 3.10 0.03 | 2.18 0.00 | 5.09 | – | * 61 Cyg B |
| 1660p363w2o0007263e000 | 165.8324 | 35.9562 | 55335 | -0.58 0.01 | -4.86 0.02 | 2.67 0.00 | 4.90 | 4.81 | [ZEH2003] RX J1103.3+3559 2 |
| 1647p075w2o0003990e000 | 164.1092 | 7.0070 | 55346 | -3.85 0.03 | -2.71 0.06 | 5.58 0.00 | 4.71 | 4.72 | Wolf 359 |
| 3323m576w2o0029882e000 | 331.0651 | -56.7899 | 55318 | 3.98 0.02 | -2.51 0.00 | 9.48 0.00 | 4.71 | – | * eps Ind B |
| 1664p439w2o0006206e001 | 166.3509 | 43.5296 | 55524 | -4.50 0.04 | 1.02 0.06 | 4.25 0.00 | 4.61 | – | BD+44 2051 |
| 3295m576w2o0026103e002 | 330.8694 | -56.7959 | 56779 | 3.72 0.09 | -2.70 0.06 | 1.70 0.00 | 4.59 | – | * eps Ind |
| 0640m076w2o0011226e002 | 63.8308 | -7.6690 | 56705 | -2.07 0.06 | -3.55 0.06 | 5.43 0.00 | 4.11 | 4.08 | * omi02 Eri C |
| 2042p030w2o0024100e000 | 204.1223 | 3.6764 | 55210 | -3.76 0.01 | -1.13 0.01 | 12.54 0.01 | 3.92 | 3.90 | Wolf 489 |
| 0640m076w2o0011262e002 | 63.8090 | -7.6663 | 56705 | -2.01 0.09 | -3.37 0.03 | 1.98 0.00 | 3.92 | 4.09 | * omi02 Eri b |
| 0178p545w2o0029817e000 | 17.0849 | 54.9158 | 55222 | 3.46 0.03 | -1.55 0.02 | 2.94 0.00 | 3.79 | – | * mu. Cas |
| 1124p045w2o0033512e000 | 111.8539 | 5.2153 | 55294 | 0.46 0.04 | -3.71 0.01 | 4.30 0.00 | 3.74 | 3.74 | BD+05 1668 |
| 2279m167w2o0018036e000 | 227.5511 | -16.4729 | 55237 | -1.02 0.01 | -3.54 0.02 | 7.12 0.00 | 3.69 | 3.68 | HD 134440 |
| 2279m167w2o0019629e000 | 227.5516 | -16.3893 | 55237 | -1.00 0.02 | -3.54 0.02 | 6.96 0.00 | 3.68 | 3.68 | HD 134439 |
| 2671p832w2o0006516e000 | 267.4266 | 82.7833 | 55273 | -1.43 0.01 | 3.30 0.01 | 13.44 0.01 | 3.60 | – | EGGR 199 |
| 2208m197w2o0002522e000 | 221.0761 | -20.3292 | 55230 | -2.92 0.00 | -1.97 0.01 | 11.22 0.00 | 3.52 | 3.51 | SSSPM J1444-2019 |
| 3193m394w2o0023066e001 | 319.3007 | -38.8709 | 55497 | -3.05 0.07 | -1.08 0.03 | 2.55 0.00 | 3.24 | – | V* AX Mic |
| 1744p666w2o0027744e000 | 175.0866 | 67.2500 | 55317 | 0.25 0.01 | -3.17 0.00 | 8.43 0.00 | 3.18 | 3.17 | Ross 451 |
| 2492m682w2o0003788e000 | 249.9201 | -68.7940 | 55265 | 0.58 0.02 | -3.12 0.01 | 13.57 0.02 | 3.17 | – | WISE J163940.83-684738.6 |
| 0041p439w2o0017537e002 | 4.6074 | 44.0241 | 55571 | 3.11 0.07 | 0.52 0.03 | 3.41 0.00 | 3.15 | – | HD 1326c |
| 0505m425w2o0002411e000 | 49.9934 | -43.0678 | 55211 | 3.04 0.05 | 0.72 0.03 | 2.07 0.00 | 3.12 | – | * e Eri |
| 0540m122w2o0024847e000 | 54.5695 | -11.4946 | 55230 | 1.47 0.00 | -2.71 0.01 | 8.54 0.00 | 3.08 | 3.07 | Ross 578 |
| 1683m258w2o0004191e000 | 168.7034 | -26.3077 | 55364 | -3.04 0.01 | -0.38 0.01 | 12.30 0.01 | 3.06 | – | 2MASS J11145133-2618235 |
| 0121p060w2o0001682e000 | 12.2948 | 5.3807 | 55383 | 1.24 0.01 | -2.74 0.00 | 11.36 0.00 | 3.00 | 2.98 | Wolf 28 |
| 2897m455w2o0014661e000 | 290.2025 | -45.5665 | 55291 | 0.71 0.03 | -2.89 0.02 | 6.52 0.00 | 2.97 | 2.97 | L 347-14 |
| 1712p651w2o0031078e000 | 169.9995 | 65.8469 | 55317 | -2.95 0.02 | 0.14 0.02 | 5.17 0.00 | 2.95 | – | V* SZ UMa |
| 1627m531w2o0012782e000 | 162.3154 | -53.3185 | 55205 | -2.79 0.03 | 0.40 0.03 | 7.35 0.00 | 2.82 | – | Luhman 16 |
| 1747m652w2o0030973e000 | 176.4464 | -64.8425 | 55224 | 2.65 0.01 | -0.33 0.01 | 11.09 0.00 | 2.67 | – | LAWD 37 |
| 0812p832w2o0005622e000 | 83.2730 | 82.7745 | 55268 | 2.03 0.00 | -1.67 0.01 | 13.25 0.01 | 2.63 | – | 2MASS J05325346+8246465 |
| 2807m637w2o0010334e000 | 281.2887 | -63.9615 | 55282 | 2.55 0.02 | 0.60 0.01 | 7.85 0.00 | 2.62 | – | SCR J1845-6357 |
| 0438p030w2o0002613e000 | 43.5394 | 2.3996 | 55223 | 2.60 0.01 | 0.21 0.01 | 12.72 0.01 | 2.61 | – | 2MASS J02540788+0223563 |
| 3432m061w2o0000958e000 | 343.4797 | -6.7837 | 55350 | 2.49 0.01 | -0.72 0.01 | 13.25 0.02 | 2.59 | 2.58 | EGGR 453 |
$1$$1$footnotetext: Unique ID corresponds to unwise_detid field of the first detection of the object in unTimely epochal catalogues.
$2$$2$footnotetext: Coordinates and epoch here correspond to predicted position of the object at the time of first detection. $3$$3$footnotetext: W2 magnitudes and errors are computed as simple averages of corresponding epochal values, without any additional filtering.
$4$$4$footnotetext: Proper motion of the closest Gaia DR3 match.
$5$$5$footnotetext: First identifier for a closest matching object from Simbad. See the text for details on how exactly the association is performed.
| CWISE | RA | Dec | pm | W2 | M_W2 | SpT | Distance | V | ||
| J2016.0 | J2016.0 | ′′/year | pc | km/s | ||||||
| 1 | J053046.20+440849.2 | 82.6924 | 44.1473 | 1.22 0.03 | 14.25 0.02 | 12.90 0.41 | T0–T81212121212121212S | 19 4 | 107 20 | |
| 2 | J011552.59-454814.1 | 18.9688 | -45.8040 | 1.02 0.07 | 15.58 0.04 | 10.36 1.02 | M8–L11212121212121212S | 111 52 | 535 251 | |
| 3 | J062316.19+071505.6 | 95.8174 | 7.2516 | 0.95 0.03 | 14.86 0.03 | 12.68 0.40 | T4–T61313131313131313S | 27 5 | 122 23 | |
| 4 | J124909.08+362116.0 | 192.2880 | 36.3548 | 0.91 0.07 | 15.59 0.06 | 9.82 0.67 | L11414141414141414S | 142 44 | 617 191 | (1) |
| 5 | J171342.26-443112.5 | 258.4262 | -44.5201 | 0.90 0.02 | 13.09 0.02 | 9.82 0.64 | L0–L61212121212121212S | 45 13 | 193 57 | |
| 6 | J105018.24-683054.9 | 162.5774 | -68.5153 | 0.87 0.07 | 15.15 0.02 | 12.99 0.54 | T6–T71313131313131313S | 27 7 | 112 28 | |
| 7 | J214706.01-540407.6 | 326.7740 | -54.0688 | 0.87 0.05 | 15.21 0.03 | 13.53 0.41 | T4–T61212121212121212S | 22 4 | 89 17 | |
| 8 | J011816.64-723917.9 | 19.5654 | -72.6550 | 0.85 0.04 | 15.39 0.03 | 12.63 0.32 | T4–T61313131313131313S | 36 5 | 144 21 | |
| 9 | J234627.27+665732.6 | 356.6121 | 66.9591 | 0.84 0.04 | 14.80 0.03 | 9.37 0.71 | 122 40 | 483 158 | ||
| 10 | J140048.92+253759.8 | 210.2038 | 25.6333 | 0.81 0.05 | 15.55 0.04 | 13.54 0.39 | T7–T91313131313131313S | 25 5 | 97 17 | |
| 11 | J224250.43-132837.7 | 340.7102 | -13.4775 | 0.81 0.07 | 15.18 0.08 | 9.92 0.81 | M8–L01212121212121212S | 113 42 | 434 162 | |
| 12 | J065233.95+223046.8 | 103.1414 | 22.5131 | 0.81 0.03 | 11.67 0.01 | 9.23 0.70 | M9–L11212121212121212S | 31 10 | 118 38 | |
| 13 | J211250.11-052925.2 | 318.2090 | -5.4902 | 0.81 0.05 | 15.20 0.04 | 11.82 0.70 | T2–T41212121212121212S | 48 15 | 182 58 | |
| 14 | J204928.79-443146.9 | 312.3697 | -44.5298 | 0.80 0.09 | 15.33 0.03 | 14.00 0.73 | T8–Y01313131313131313S | 18 6 | 70 24 | (2) |
| 15 | J011128.00-103535.7 | 17.8664 | -10.5930 | 0.77 0.04 | 15.24 0.05 | 9.32 0.70 | L01414141414141414S | 153 50 | 557 181 | |
| 16 | J075744.48-300504.3 | 119.4353 | -30.0846 | 0.76 0.03 | 14.74 0.03 | 12.34 0.24 | T3–T51212121212121212S | 30 3 | 109 12 | |
| 17 | J141946.85-610910.3 | 214.9464 | -61.1527 | 0.75 0.03 | 11.95 0.01 | 9.22 0.58 | M5–M71212121212121212S | 35 9 | 124 33 | |
| 18 | J230252.77-131458.0 | 345.7199 | -13.2494 | 0.74 0.05 | 15.51 0.05 | 13.07 0.47 | T4–T51212121212121212S | 31 7 | 108 23 | (3) |
| 19 | J114557.19+673327.5 | 176.4898 | 67.5577 | 0.72 0.05 | 15.88 0.05 | 11.74 0.53 | T2–T41313131313131313S | 67 16 | 230 56 | |
| 20 | J021059.50-395414.2 | 32.7480 | -39.9040 | 0.70 0.04 | 15.60 0.04 | 9.78 0.66 | M8–L41212121212121212S | 146 44 | 487 148 | |
| 21 | J045456.23-742522.6 | 73.7352 | -74.4229 | 0.70 0.02 | 15.48 0.03 | 13.41 0.92 | T7–T81313131313131313S | 26 11 | 86 36 | |
| 22 | J104222.34-834921.2 | 160.6012 | -83.8227 | 0.69 0.03 | 15.03 0.03 | 9.54 0.75 | M7–M91212121212121212S | 125 43 | 411 142 | (4) |
| 23 | J024641.43-634329.2 | 41.6714 | -63.7248 | 0.68 0.03 | 15.21 0.03 | 12.63 0.32 | T4–T51212121212121212S | 33 5 | 106 16 | |
| 24 | J110048.87-554006.8 | 165.2038 | -55.6685 | 0.68 0.02 | 14.15 0.02 | 9.54 0.76 | M6–L11212121212121212S | 83 29 | 268 93 | (5) |
| 25 | J095855.00+540305.4 | 149.7298 | 54.0515 | 0.68 0.03 | 15.27 0.03 | 12.83 0.28 | T5–T61313131313131313S | 31 4 | 98 13 | |
| 26 | J222721.75+325119.5 | 336.8402 | 32.8554 | 0.67 0.03 | 14.97 0.03 | 11.18 0.79 | T2–T31212121212121212S | 57 21 | 182 66 | |
| 27 | J060302.31+321532.1 | 90.7595 | 32.2589 | 0.66 0.04 | 15.27 0.04 | 9.38 0.74 | M7–M91212121212121212S | 151 51 | 473 161 | (6) |
| 28 | J201029.81+015100.3 | 302.6243 | 1.8501 | 0.66 0.07 | 15.19 0.03 | 9.48 0.76 | M9–L51212121212121212S | 138 49 | 434 153 | |
| 29 | J210819.58+182318.3 | 317.0814 | 18.3884 | 0.65 0.02 | 14.31 0.02 | 9.86 0.73 | L0–L71212121212121212S | 78 26 | 239 80 | |
| 30 | J092710.37-474155.5 | 141.7937 | -47.6986 | 0.64 0.04 | 14.69 0.03 | 13.22 0.34 | T6–T81313131313131313S | 20 3 | 59 9 | |
| 31 | J194428.05+053344.8 | 296.1168 | 5.5622 | 0.63 0.02 | 13.82 0.04 | 9.85 0.76 | M9–T31212121212121212S | 62 22 | 187 66 | (7) |
| 32 | J220134.02-041609.5 | 330.3919 | -4.2694 | 0.62 0.04 | 15.21 0.04 | 12.91 0.40 | T4–T61212121212121212S | 29 5 | 85 16 | (8) |
| 33 | J102332.99+221527.6 | 155.8875 | 22.2577 | 0.61 0.06 | 15.23 0.04 | 12.75 0.31 | T4–T61313131313131313S | 31 5 | 91 13 | |
| 34 | J141403.72-171531.1 | 213.5154 | -17.2588 | 0.61 0.03 | 15.33 0.04 | 12.68 0.38 | T4–T61212121212121212S | 34 6 | 98 17 | |
| 35 | J184526.47+754829.7 | 281.3607 | 75.8084 | 0.60 0.02 | 15.12 0.02 | 12.47 0.39 | T4–T51313131313131313S | 34 6 | 97 17 | |
| 36 | J105716.29-675728.4 | 164.3176 | -67.9571 | 0.59 0.03 | 13.86 0.05 | K21414141414141414S | (9) | |||
| 37 | J010803.06-204741.1 | 17.0127 | -20.7948 | 0.59 0.03 | 14.49 0.02 | 9.74 0.65 | M9–L11212121212121212S | 89 27 | 248 74 | |
| 38 | J065919.42+161551.8 | 104.8309 | 16.2643 | 0.58 0.04 | 15.32 0.04 | 11.66 0.57 | T0–T31212121212121212S | 54 14 | 148 39 | |
| 39 | J140808.52-491759.4 | 212.0358 | -49.3000 | 0.58 0.03 | 14.54 0.02 | 12.73 0.35 | T51313131313131313S | 23 4 | 63 10 | |
| 40 | J035532.23+474358.1 | 58.8838 | 47.7328 | 0.56 0.03 | 14.30 0.02 | 13.05 0.52 | T5–T71313131313131313S | 18 4 | 48 11 | (10) |
| 41 | J093016.98+604631.2 | 142.5714 | 60.7754 | 0.55 0.03 | 14.92 0.02 | 13.00 0.72 | T5–T71313131313131313S | 24 8 | 63 21 | (11) |
| 42 | J062158.80+220048.5 | 95.4951 | 22.0131 | 0.52 0.03 | 14.02 0.02 | 9.85 0.64 | L1–L61212121212121212S | 68 20 | 170 50 |
$1$$1$footnotetext: Published in Burgasser et al. (2024) as an sdL1 subdwarf with 0.8840.005′′/year proper motion at 1258 pc distance having 10310 km/s radial velocity.
$2$$2$footnotetext: Published in Tinney et al. (2018) with estimated T7.3 spectral type, and included in UltraCoolSheet (Best et al., 2024).
$3$$3$footnotetext: Published in Brooks et al. (2024) with L1 spectral type and 142 pc distance. See Section 3.2 for more detailed discussion.
$4$$4$footnotetext: Published in Brooks et al. (2024) with M8 spectral type and 126 pc distance.
$5$$5$footnotetext: Published in Brooks et al. (2024) with M9 spectral type and 77 pc distance.
$6$$6$footnotetext: Published in Brooks et al. (2024) with M8 spectral type and 138 pc distance.
$7$$7$footnotetext: Published in Brooks et al. (2024) with M4 spectral type and 509 pc distance.
$8$$8$footnotetext: Published in Brooks et al. (2024) with L1 spectral type and 174 pc distance. See Section 3.2 for more detailed discussion.
$9$$9$footnotetext: CatWISE2020 association unreliable. WISE magnitdes computed from unTimely measurements. Optical and IR colors consistent with early K-type star.
$10$$10$footnotetext: Published in Smith et al. (2014) with estimated T6 spectral type at approximately 19 pc distance, and included in UltraCoolSheet (Best et al., 2024).
$11$$11$footnotetext: Published in Brooks et al. (2024) with T6 spectral type and 23 pc distance.
pectral type estimated from the position in (K-W2) - (J-W2) plane.
pectral type estimated from (W1-W2) color.
pectral type estimated from VOSA (Bayo et al., 2008) template fitting of multi-wavelength measurements.
| CWISE | cW1 | cW2 | 2MASS | J | H | K | |
| 1 | J053046.20+440849.2 | 15.990.03 | 14.250.02 | 05304606+4409108 | 16.380.1 | 16.140.18 | 17.12 |
| 2 | J011552.59-454814.1 | 16.00.02 | 15.580.04 | ||||
| 3 | J062316.19+071505.6 | 16.480.04 | 14.860.03 | ||||
| 4 | J124909.08+362116.0 | 15.920.03 | 15.590.06 | ||||
[…]

### v2

@@ 3 Discussion
We successfully detected 21,885 moving objects in Section 2.1, including the largest proper motion brown dwarf, WISEA J085510.74-071442.5(Luhman, 2014), as well as the fastest known star, Barnard’s Star, and the majority of known high proper motion ultracool dwarfs from UltracoolSheet (Best et al., 2024, 2021, 2024). However, we did not detect anything moving faster than that, despite the algorithm specifically being tailored to be sensitive to even the largest possible proper motions (see Figure 4).
Our algorithm facilitated the identification of 21,885 confirmed moving objects from 59,322 candidates through visual inspection. This approach demonstrated a notable advancement over prior methods. For instance, the image-based neural network SMDET, applied to unWISE coadds, identified 1,730 real proper motion objects from 11,900 inspected candidates (Brooks et al., 2024). Additionally, our algorithm uncovered 323 moving objects without SIMBAD and Gaia associations – nearly three times more than the 118 reported in a similar range of proper motions by Brooks et al. (2024) – while successfully recovering 38 objects from their sample.
The process of visually inspecting initial candidates and identifying moving objects based on their visual signatures (see Figure 1) closely resembles the approach used by the Backyard Worlds: Planet 9 citizen science project (kuchner_2017), which involves tens of thousands of participants and has been ongoing for seven years. However, by leveraging a motion detection algorithm to generate track candidates, applying a machine learning-based artifact rejection routine, and utilizing a compact, Jupyter-based dashboard, our visual inspection process is significantly more fast and efficient.
Backyard Worlds: Cool Neighbors (Humphreys et al., 2020) citizen science project refines the original by also using an input list of positions – in their case, selected from CatWISE2020(Marocco et al., 2021) catalogue using machine learning classifier tailored specifically for finding new T and L brown dwarfs, with the expected efficiency of about 0.5% due to large amount of spurious high proper motion entries polluting the catalogue. In contrast to them, our method is not restricted to finding just brown dwarfs, and is better suited for detecting high proper motion objects which is difficult for CatWISE2020(Marocco et al., 2021). Moreover, among 258 new moving objects we detected three that are lacking CatWISE2020 associations, proving that our approach is also complementary to the ones based on that catalogue like Humphreys et al. (2020) or Kota et al. (2022).
Figure 9 shows sky positions of the newly detected objects in Galactic coordinates. It is consistent with isotropic distribution expected for nearby Galactic objects. We may qualitatively assess the distances to them by regressing the absolute magnitudes of objects with known parallaxes from 20-pc census of stars and brown dwarfs (Kirkpatrick et al., 2024) versus their spectral types, and then applying the same dependency (along with corresponding scatter) for best spectral type estimates from our sample, as described in Section 2.3. It then allows (if we ignore the reddening which is reasonable for such nearby objects) estimating the distance (“photometric distance”) using observed magnitudes with formal accuracy typically better than 20%, as well as estimating projected velocities with comparable accuracy, as shown in Figure 11. Comparison of these photometric distances with astrometric ones for the objects we detected in Section 2.3 that lack SIMBAD associations but do have Gaia DR3 or DR2 counterpart is shown in left panel of Figure 10. While Gaia detected objects all have spectral types earlier than M9, we may still see that photometric distances are, within the errors, compatible with the parallaxes, for most of the colder objects. Starting with approximately M3 spectral type and earlier, the estimation becomes biased and significantly (by about factor of two) overestimates the distance for late K types. We cannot easily guess the reason for this behaviour, but it may explain the anomalously high distances and projected velocities for two hottest objects in our actual new objects sample shown in Figure 11. The rest of the objects there should not be affected by this problem, and we may consider their photometric distances reliable. On the other hand, the agreement between the effective temperatures estimated in Section 2.3 with the ones derived using Gaia DR3 GSP-Phot pipeline (Andrae et al., 2023) is sufficiently good up to at least TK, or K6 spectral type (see right panel of Figure 10.
Among 258 new objects we detected all except 6 are compatible with being ultracool dwarfs colder than M7 type. At least 62 are consistent with being brown dwarfs of types later than T0, i.e. having K or . Among them, 33 have effective temperatures estimated using BT-Settl spectral grid, and thus may be considered most reliable new T brown dwarf candidates that require further spectroscopic study for their confirmation. They all have estimated distances of 20–40 pc, with the closest one – CWISE J154802.28+565802.2 – being at a pc distance and with T8 estimated spectral type. We did not detect any new hypervelocity star (Brown, 2015) (two objects with formal projected velocity exceeding 1000 km/s in Figure 11 are most probably due to overestimated distance, see above), and no brown dwarf candidate with projected velocity significantly greater than km/s.
[…]